Average Rate of Return Method, Net Present Value Method, and Analysis for a service company

The capital investment committee of Arches Landscaping Company is considering two capital investments. The estimated income from operations and net cash flows from each investment are as follows:

Front-End Loader Greenhouse
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $59,400 $195,000 $125,000 $312,000
2 59,400 195,000 95,000 263,000
3 59,400 195,000 48,000 185,000
4 59,400 195,000 21,000 127,000
5 59,400 195,000 8,000 88,000
Total $297,000 $975,000 $297,000 $975,000

Each project requires an investment of $540,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Front-End Loader %
Greenhouse %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Front-End Loader Greenhouse
Present value of net cash flow $ $
Amount to be invested $ $
Net present value $ $

2. Prepare a brief report for the capital investment committee, advising it on the relative merits of the two investments.

The front-end loader has a   net present value because the cash flows occur   in time compared to the greenhouse. Thus, if only one of the two projects can be accepted, the   would be the more attractive.
Answer #1
Average net income 59400 =297000/5
Average investment 270000 =540000/2
1a
Average Rate of Return
Front-End Loader 22% =59400/270000
Greenhouse 22% =59400/270000
1b
Front-End Loader Greenhouse
Present value of net cash flow 739050 781170
Amount to be invested 540000 540000
Net present value 199050 241170
2
The front-end loader has a smaller net present value because the cash flows occur later in time compared to the greenhouse. Thus, if only one of the two projects can be accepted, the Greenhouse would be the more attractive.
Workings:
Front-End Loader Greenhouse
Cash flows PV factor Present value Cash flows PV factor Present value
1 195000 0.909 177255 312000 0.909 283608
2 195000 0.826 161070 263000 0.826 217238
3 195000 0.751 146445 185000 0.751 138935
4 195000 0.683 133185 127000 0.683 86741
5 195000 0.621 121095 88000 0.621 54648
Total 739050 Total 781170